# PC-Gym: Benchmark Environments for Process Control Problems

### v1

@@ Abstract
PC-Gym is an open-source tool designed to facilitate the development and evaluation of reinforcement learning (RL) algorithms for chemical process control problems. It provides a suite of environments that model a range of chemical processes, incorporating nonlinear dynamics, process disturbances, and constraints. Key features include flexible constraint handling mechanisms, customizable disturbance generation, and modular reward function design. The framework enables benchmarking state-of-the-art RL algorithms against a nonlinear Model Predictive Control (NMPC) oracle across various process control scenarios. Case studies demonstrate PC-Gym’s effectiveness in evaluating RL approaches for the control of various chemical engineering systems such as a continuously stirred tank reactor, multistage extraction process, and crystallization reactor. The framework’s ability to incorporate realistic disturbances and constraints allows for robust testing of control strategies. Results highlight the performance gaps between RL algorithms and NMPC oracles, demonstrating the utility of PC-Gym for algorithm benchmarking and suggesting areas for improvement in RL-based process control. By offering a standardized platform for developing and assessing RL-based control strategies, PC-Gym aims to accelerate research at the intersection of machine learning and process systems engineering. It bridges the gap between theoretical advancements in RL and practical applications in industrial process control, providing researchers and practitioners with a valuable tool for exploring data-driven control solutions for complex chemical processes.

### v2

@@ Abstract
PC-Gym is an open-source tool for developing and evaluating reinforcement learning (RL) algorithms in chemical process control. It features environments that simulate various chemical processes, incorporating nonlinear dynamics, disturbances, and constraints. The tool includes customizable constraint handling, disturbance generation, and reward function design. The tool enables comparison of RL algorithms against nonlinear Model Predictive Control (NMPC) across different scenarios. Case studies demonstrate its effectiveness in evaluating RL approaches for systems like continuously stirred tank reactors, multistage extraction processes, and crystallization reactors. The results reveal performance gaps between RL algorithms and NMPC oracles, highlighting areas for improvement. By providing a standardized platform, PC-Gym aims to accelerate research at the intersection of machine learning and process systems engineering. It connects theoretical RL advances with practical industrial process control applications, offering researchers a tool for exploring data-driven control solutions.